OpenStudy (anonymous):

what is an equation in standard form that "has a slope of (2) and a y intercept of (3)

OpenStudy (anonymous):

do you mean slope intercept form?

OpenStudy (anonymous):

idk this is what the equation

OpenStudy (anonymous):

said

OpenStudy (polaris_s0i):

y = 2x + 3 if you have 2 points on the line: \[p_1(x_1,y_1),\space p_2(x_2,y_2)\] the form is: \[y = \frac{\Delta y}{\Delta x}*(x - xint) + yint\] or \[y = \frac{y_2 - y_1}{x_2 - x_1}*(x - xint) + yint\] more commonly people use m for slope so:and b for the y intercept. \[y = mx + b\]